OpenStudy (anonymous):

Evalute w/o using a calulator! csc(-420) Please explain & show work because i'm very new to this!

OpenStudy (dumbcow):

find reference angle take away 360 since that is one time around unit circle csc(-60) you are 60 degrees away from axis in 4th quandrant (csc is neg) csc x = 1/sinx --> sin(60) = sqrt3/2 --> csc(60) = 2/sqrt3 finally make it neg because you are in 4th quadrant -2/sqrt3

OpenStudy (phi):

no wait. csc(x) = 1/sin(x) and sin(-60)= -sqrt(3)/2 so -2/sqrt(3) or -2 sqrt(3)/3 sure and steady (like dumbcow) is the best way...

OpenStudy (phi):

But you should memorize cos(60) = 1/2 sin(60)= sqrt(3)/2 tan(60)= sqrt(3)

OpenStudy (anonymous):

thank you:)